OpenStudy (anonymous):

Problem Set 5 Incorrect Solution?

OpenStudy (anonymous):

Can anyone verify that Integration Problem Set 5, problem 3A-2E provides an incorrect solution? Using substitution where: \[u = 8-2x^{2}, du = -4xdx\] … the solution then goes on to assume \[\int\limits_{}^{-} xdx/\sqrt{8-2x^{2}} = \int\limits_{-}^{-}u^{1/2}(-1/4)du\]… but shouldn't that be… a -1/2 power instead, resulting in: \[=\int\limits_{-}^{-}u^{-1/2}(-1/4)du\]? The final solution I get: \[-2\sqrt{8-2x^{2}}\]

OpenStudy (anonymous):

This is one of a number of problems for which I've noted errors in the solution. Somehow they've offered a solution for u^(1/2) rather than the problem that was actually posed, which is u^(-1/2). You may want to check your solution, though, shouldn't it be -1/2 out front?

OpenStudy (anonymous):

Good to know! I'll take a look at your answer history before re-posting any questions :) Also, you are right about my solution... I somehow thought -2/4 = -2 ha! The -1/2 out front looks better. Thanks again Creeksider!